# Exploring Knowledge Tracing in Tutor-Student Dialogues

### v1

@@ 6.2 Qualitative Analysis
We now detail findings from our qualitative analysis, where we seek to i) identify challenges in the dialogue KT task and ii) understand why LLMKT has an advantage over existing KT methods. Therefore, we examine LLMKT’s predicted KC mastery and correctness on dialogues that are in the test set from our experiments on CoMTA.
@@
We identify three primary challenges in the data that are unique to the dialogue KT task that are difficult for KT methods. First, the open-ended nature of dialogues means that student behavior is much less structured than in traditional assessment or practice settings. For example, when the tutor poses a question to the student, the student often i) does not answer directly but instead asks a clarifying question, ii) skips ahead to the final solution in the case of step-by-step problem solving, or iii) asks the tutor to give them a different problem. These student actions are inherently difficult to anticipate ahead of time since they define behavior beyond what can be captured by the KCs. As a result, there are many turns in these dialogues that are “irrelevant” to the KT task. Future work may develop methods to classify student intent and automatically identify these turns.
@@
Second, since dialogues can be short (around 4 turn pairs on average), there is little historical context to estimate student knowledge from. In many dialogues, students are proficient in one aspect of problem solving, responding to several questions correctly at first, but then struggle when the tutor moves on to a new KC. Without sufficiently long student history, it is difficult to estimate KC mastery levels for ones that are seen for the first time. This problem is partially due to existing datasets not linking students over multiple dialogues. Future work may alleviate this problem by tracking students over multiple dialogues or including additional student-level information, such as pre-test scores.
Third, the varying number of KCs in different dialogue turns can lead to numerous false negatives and false positives, due to the product in our “AND”-type correct response likelihood model (Eq. 1). Even if the predicted KC mastery levels are high, the predicted correctness probability will be low for a turn will multiple KCs. On the contrary, turns with fewer KCs are more likely to be predicted as correct. Future work can explore other more sophisticated models for turns with multiple KCs, such as DINA [12] or Boolean logic analysis [23], to identify which KCs are more important to the correctness of a response rather than treating them equally.
@@
We find that, unsurprisingly, LLMKT’s advantage over existing KT methods comes from leveraging nuanced details in the dialogue, rather than solely relying on previous correctness and KC labels. For the most part, the predicted mastery level correlates with previous turn correctness; as a result, dialogues with many correct or incorrect turns in a row are easy to predict for all KT methods. LLMKT excels in cases when it relies on textual information, rather than correctness labels, to adjust its KC mastery estimates. Table 3 shows one such example. At the first turn, the student responds incorrectly because they do not provide the answer the tutor was looking for. However, the response still shows some understanding of the topic; therefore, LLMKT increases, rather than decreases, the KC mastery prediction, despite the incorrect label for this turn. As a result, it correctly predicts that the student responds correctly to later turns while other KT methods cannot.
@@
Table 3: An example dialogue with student turn correctness labels, KC labels, and LLMKT’s predicted mastery level for each KC. Even though the student’s response at the first turn is not fully correct, it suggests that they do have some knowledge on underlying topic, causing LLMKT to increase its mastery predictions in the next steps as the student responds correctly.
Speaker Turn Text Knowledge Components - $\hat{z}_{jk}$
Student how do I multiply this? 7x $\times$ 8
Tutor Ah, you’re asking about multiplication! Let’s break it down. What do you get when you multiply $7x$ by $8$? 1) Write, read, and evaluate expressions in which letters stand for numbers. $0.6792$
Student is it possible to multiply the 8 and 7?
Tutor You’re thinking correctly! When we perform multiplication, we can rearrange the numbers in any order. So, we can first multiply 7 and 8. What’s the product of 7 and 8? 1) Apply properties of operations as strategies to multiply and divide. $0.7773$
2) Fluently multiply and divide within 100. $0.7549$
Student 56
Tutor Excellent work! Now, don’t forget we still have that $x$ from $7x$. What should we do with it? 1) Write, read, and evaluate expressions in which letters stand for numbers. $0.7773$
Student 56x?
6.3 Learning Curve Analysis
A common method for assessing the quality of KCs and KT methods is examining how well they satisfy pedagogical theory, in particular, the power law of practice [49]: as students practice a particular skill, their mastery of it will increase rapidly at first and then more slowly over time. Following prior work [47], we visualize changes in predicted mastery levels at the KC-level, examine trends in the curves, and investigate whether there is evidence of the power law of practice. We use LLMKT’s predicted masteries on the CoMTA data for all dialogues when they are in the test split of the cross-fold experiment. We use CoMTA instead of MathDial for this analysis since the latter uses LLMs to simulate students; it is unlikely for simulated students to behave like real ones and we confirmed this in a preliminary analysis.
To compute “learning curves”, for a KC $c$, we first find each time that KC occurs in the data, and we use $c^{(i)}_{j}$ to denote its $j^{\text{th}}$ occurrence in the $i^{\text{th}}$ dialogue, and use $\hat{z}^{(i)}_{j}$ to denote the corresponding predicted KC mastery level. We then subtract $\hat{z}^{(i)}_{1}$ from each $\hat{z}^{(i)}_{j}$ to track the change in predicted mastery from the first prediction, since different students may have different initial knowledge levels. We plot the average KC mastery level, i.e., $\frac{1}{N}\sum_{i=1}^{N}\hat{z}^{(i)}_{j}-\hat{z}^{(i)}_{1}$, where $N$ is the number of dialogues that have at least $j$ occurrences of $c$, and show the standard deviation using error bars.
@@ 6.3 Learning Curve Analysis
Figure 2: “Learning curves”, i.e., estimated KC mastery levels over dialogue turns, averaged over students for selected KCs in the CoMTA dataset. More common KCs tend to show slightly increasing mastery levels as the dialogue progresses, while less common ones tend to show the opposite. From left to right: “Solve linear equations and inequalities in one variable, including equations with coefficients represented by letters,” “Use variables to represent numbers and write expressions when solving a real-world or mathematical problem,” and “Apply the properties of operations to generate equivalent expressions.”
@@
We examine plots of the 15 most frequent KCs since less common KCs do not have enough data to show clear trends. Out of these, the results are mixed: 7 out of 15 KCs trend towards increasing mastery as the dialogues go on, while 8/15 KCs trend towards decreasing mastery. Figure 2 shows representative examples of each case. KCs with an upward trend typically follow the same pattern: mastery initially has high variance and features a small downward trend, but is then followed by a large and steady increase. Additionally, these KCs are more prevalent in longer dialogues and correspond to the more commonly occurring KCs in the data. These patterns can be explained by two concurrent trends in the data: First, there are several long dialogues where a student struggles with a math topic at the beginning, getting some questions correct and others incorrect, but after some time becomes more proficient and consistently responds correctly; these dialogues contribute to the steady increase in KC mastery. Second, there are several short dialogues where the student either simply knows the topic and responds correctly or struggles with the topic and responds incorrectly before the dialogue quickly ends; these dialogues contribute to the high variance and early dip in KC mastery.
@@
On the contrary, KCs with a downward trend tend to be associated with shorter dialogues and are less frequent in the dataset. These dialogues tend to feature a student repeatedly trying to solve a problem without any success, only to end the dialogue immediately after getting to the correct solution. KC mastery can also drop significantly at the beginning of the dialogue once the student starts to respond incorrectly. Since there are few dialogues to average over, these downward trends can dominate the learning curves; the most common KC with a downward trend only appears in 8 dialogues, which is less than most KCs with upward trends.
@@
Overall, when we take into account the small amount of data, the learning curve trends seem to mostly resemble the power law of practice; when there are sufficient turns in a dialogue, the student’s predicted mastery eventually increases before starting to level out. The main reason these trends are not more prevalent is likely due to the dialogues typically ending after only a few turns. This setting is different than the problem-solving setup studied in existing KT methods, where students practice a KC repeatedly by solving different problems, even after mostly mastering it. Again, we call upon researchers who have access to data with longer dialogues to conduct more extensive analyses.
@@ 7 Annotation Quality Experiments
7.1 Human Evaluation Setup and Results
To quantify GPT-4o’s accuracy on the correctness and KC labeling tasks, we conduct an IRB-approved human evaluation using expert human annotators. We hired 3 former math teachers at \$40 US/hour, each rating the same set of 30 randomly selected dialogues from CoMTA with GPT-4o’s labels, totaling 166 turn pairs. For each turn in each dialogue, we ask annotators to score GPT-4o’s label i) on correctness, as 1 if they think the label is accurate and 0 otherwise, and ii) on the standards selected by GPT-4o on a 1-to-4 scale with 4 as the highest. We provide the exact instructions given to annotators in an online Supplementary Material [^4] https://osf.io/wrkej?view_only=f32472a6560649e692a4e2ba603ef52c. For both correctness and KCs, we report the average human-assigned score and two inter-rater reliability (IRR) metrics: i) the portion of times that the labels of all 3 annotators are exactly the same (Overlap) and ii) the Krippendorff’s $\mathbf{\alpha}$ coefficient [5]. We also compute the accuracy of correctness predictions at the final turn using CoMTA’s ground truth labels for both GPT-4o and our annotators.
Table 4 shows the human evaluation results. We see that our annotators give GPT-4o very high scores for correctness and moderate-to-high scores for KCs, indicating that GPT-4o is able to competently perform these tasks. IRR is volatile, with high overlap and low $\alpha$ on correctness, and low overlap but moderate $\alpha$ on KCs. However, the low $\alpha$ on correctness is expected given the high class imbalance of the data (more than 90% of labels are 1), and the low overlap on KCs is expected since there are 4 classes and exact ratings can be subjective. GPT-4o’s accuracy on final turn correctness is lower than the annotator score might suggest, though this is expected since the final turns are much harder to compute correctness on without subsequent tutor turns to help indicate the label. As reference, the CoMTA tech report [35] reports final turn correctness evaluation with GPT-4o at similar accuracy (78.3). Finally, given that the average human annotator’s final turn correctness accuracy is only slightly higher than GPT-4o’s, we can conclude that the correctness labeling task is a challenging one where GPT-4o achieves close to human-level performance.
@@ 7.2 Error Analysis
To understand when GPT-4o’s errors in labeling correctness and KCs happen, we qualitatively analyze its annotations. When labeling correctness, most of GPT-4o’s errors are on the final turn, especially in questions that require some numerical calculations to verify correctness, such as when the student provides the numerical result of a sum or product calculation. This is a known limitation in GPT-4 [45] and can possibly be improved through external tools such as calculators. The other common error tends to be when GPT-4o should label a turn as “na” but does not, e.g., when the tutor asks the student whether they know how to use a function on their calculator. GPT-4o also sometimes labels a turn as correct when the tutor asks a student if they know a concept and the student indicates they do not, showing that GPT-4o can be confused by the tutor’s intentions.
@@
8 Conclusions and Future Work
In this paper, we introduced the task of dialogue KT with the goal of estimating student knowledge from tutor-student dialogues. We used GPT-4o to automatically annotate student response correctness and relevant KCs in the dialogue turns, and showed via an expert human evaluation that these annotations are largely accurate. We introduced a new LLM-based KT method that leverages the textual content in dialogues, and showed that LLMKT i) significantly outperforms existing methods on dialogue KT and ii) works with limited training data. We also showed that our models use the textual content in dialogues to capture student knowledge in meaningful ways.
Considering that this is the first work to attempt KT in dialogues, we acknowledge the limitations of our work but hope it will open up numerous avenues for future work. First, advanced prompting techniques can increase accuracy on the dialogue annotation tasks, and more sophisticated LLM-based methods can improve performance on the KT task. Second, the scope of dialogue KT can be expanded upon: in this work, we only focus on math dialogues between a single student and tutor, but future work can extend to other subjects and dialogues involving multiple students. Third, there are many potential applications of dialogue KT that can be explored. Future work should investigate whether estimated knowledge states are useful for personalization, providing teacher support, or even used to train AI-based tutors by powering simulated student agents. Overall, we hope that this initial investigation can broaden the scope of student modeling to conversational settings, which have become of significant interest lately.
9 Acknowledgement
The authors would like to thank Kristen Dicerbo and the Khan Academy team for guidelines on how to use the CoMTA dataset. We would also like to thank Nigel Fernandez and Zichao Wang for helpful discussions around this work. This work is partially supported by Schmidt Futures under the learning engineering virtual institute (LEVI) initiative and the NSF under grants 2118706, 2237676, and 2341948.

### v2

@@ 6.2 Qualitative Analysis: Case Study
We now detail findings from our qualitative analysis, where we seek to i) identify challenges in the dialogueKT task, ii) understand why LLMKT has an advantage over existing KT methods, and most importantly, iii) show how LLMKT can be useful to tutors in practice. Therefore, we perform a case study using a dialogue in the test set of CoMTA.
@@
We identify two primary challenges in the data that are unique to the dialogueKT task and are difficult for prior KT methods. First, the open-ended nature of dialogues means that student behavior is much less structured than in traditional assessment or practice settings. For example, when the tutor poses a question to the student, the student often i) does not answer directly but instead asks a clarifying question, ii) skips ahead to the final solution in the case of step-by-step problem solving, or iii) asks the tutor to give them a different problem. These student actions are inherently difficult to anticipate ahead of time since they define behavior beyond what can be captured by the KCs. As a result, there are many turns in these dialogues that are “irrelevant” to the KT task.
@@
Second, since dialogues can be short (around 4 turn pairs on average in CoMTA), there can be little historical context to estimate student knowledge from. In many dialogues, students are proficient in one aspect of problem solving, responding to several questions correctly at first, but then struggle when the tutor moves on to a new KC. Without sufficiently long student history, it is difficult to estimate KC mastery levels for ones that are seen for the first time. This problem is partially due to existing datasets not linking students over multiple dialogues.
@@
We find that, unsurprisingly, LLMKT’s advantage over existing KT methods comes from leveraging nuanced details in the dialogue, rather than solely relying on previous correctness and KC labels. For the most part, changes in predicted mastery levels correlate with previous turn correctness; as a result, dialogues with many correct or incorrect turns in a row are easy to predict for all KT methods. LLMKT excels in cases when it relies on textual information, rather than correctness labels, to adjust its KC mastery estimates. Table 3 shows one such example: At the first turn, the tutor asks the student for relevant information regarding the problem, and the student responds correctly by applying the Pythagorean Theorem, which the model accurately predicts. However, after the tutor asks the student to find the coordinates of Y, the student incorrectly wants to divide the segment by 3 rather than 4, which the model also accurately predicts. The trained model seems to have estimated that the second tutor question is more difficult than the first, using the exact textual content of the tutor question, even though they have overlapping KC labels. This correct prediction is not possible without the question parsing ability enabled by using an LLM as the underlying KT model.
@@
6.3 “Learning Curve” Analysis
@@
Figure 2: Knowledge change curves, i.e., estimated KC mastery level changes across dialogue turns, averaged over students for selected KCs in the CoMTA dataset. More common KCs in longer dialogues tend to show increasing mastery levels as the dialogue progresses, while less common KCs tend to show the opposite. KCs shown: “Solve linear equations and inequalities in one variable,” “Understand that polynomials form a system analogous to the integers,” and “Apply the properties of operations to generate equivalent expressions.”
A common method for assessing the quality of KCs and KT methods is examining how well they match pedagogical or cognitive theory, and in particular, the power law of practice [54]: as students practice a particular skill, their mastery will increase rapidly at first and then more slowly over time. Following prior work [52], we visualize changes in predicted mastery levels at the KC-level, examine trends in the curves, and investigate whether there is evidence of the power law of practice. We use LLMKT’s predicted masteries on CoMTA for all dialogues when they are in the test split of the cross-fold experiment. We use CoMTA instead of MathDial for this analysis since the latter uses LLMs to simulate students; it is unlikely for simulated students to behave like real ones, which we found in a preliminary analysis.
For a KC $c$, we first find each time that KC occurs in the data, use $c^{(i)}_{j}$ to denote its $j^{\text{th}}$ occurrence in the $i^{\text{th}}$ dialogue, and use $\hat{z}^{(i)}_{j}$ to denote the corresponding predicted KC mastery level. We then subtract $\hat{z}^{(i)}_{1}$ from each $\hat{z}^{(i)}_{j}$ to track the change in predicted mastery from the first prediction, since different students may have different initial knowledge levels. We plot the average KC mastery level, i.e., $\frac{1}{N}\sum_{i=1}^{N}\hat{z}^{(i)}_{j}-\hat{z}^{(i)}_{1}$, where $N$ is the number of dialogues that have at least $j$ turns that have $c$, and show the standard deviation using error bars.
@@
We examine plots of the 15 most frequent KCs since less common KCs do not have enough data to show clear trends. Out of these, the results are mixed: 5 out of 15 KCs trend towards increasing mastery as the dialogues go on, 5 KCs trend towards decreasing mastery, and 5 KCs do not change significantly in mastery. Figure 2 shows a representative example of each case. KCs with an upward trend typically have few changes in their masteries for the first few turns, as well as relatively high variance. Additionally, these KCs are more prevalent in longer dialogues and correspond to the more commonly occurring KCs in the data. These patterns can be explained by two concurrent trends in the data: First, there are several long dialogues where a student struggles with a math topic at the beginning, getting some questions correct and others incorrect, but after some time becomes more proficient and consistently responds correctly; these dialogues contribute to the steady increase in KC mastery. Second, there are several short dialogues where the student either simply knows the topic and consistently responds correctly, or struggles with the topic and responds incorrectly; these dialogues contribute to the high variance and the minimal change in early turns.
@@
On the contrary, KCs with a downward trend or no trend at all tend to be associated with shorter dialogues and occur in fewer dialogues. These dialogues tend to feature a student repeatedly trying to solve a problem without any success, with the dialogue often ending immediately after a correct solution is reached. KC mastery can also drop significantly at the beginning of the dialogue once the student starts to respond incorrectly. Since there are few dialogues to average over, these downward trends can dominate the knowledge change curves; the most common KC with a downward trend only appears in 8 dialogues, which is less than most KCs with upward trends.
@@
Overall, when we take into account the small amount of data and how dialogues in the CoMTA dataset are truncated, the knowledge change curve trends seem to mostly resemble the power law of practice; when there are sufficient turns in a dialogue, the student’s predicted mastery increases before leveling out eventually. The main reason these trends are not more prevalent is likely due to these dialogues typically ending after only a few turns. This setting is different than the problem-solving setup mostly used in past KT research, where students practice a KC repeatedly by solving different problems, even after mostly mastering it. Again, we call upon researchers who have access to data with longer dialogues to conduct more extensive analyses.
@@ 7 Annotation Quality Experiments
7.1 Human Evaluation
To quantify GPT-4o’s accuracy on the correctness and KC labeling tasks, we conduct an IRB-approved human evaluation using expert human annotators. We hired 3 former math teachers at \$40 US/hour, each rating the same set of 30 randomly selected dialogues from CoMTA with GPT-4o’s labels, totaling 166 turn pairs. For each turn in each dialogue, we ask annotators to score GPT-4o’s label i) on correctness, as 1 if they think the label is accurate and 0 otherwise, and ii) on the standards selected by GPT-4o on a 1-to-4 scale with 4 as the highest. We provide the exact instructions given to annotators in the online Supplementary Material [^4] https://osf.io/wrkej?view_only=f32472a6560649e692a4e2ba603ef52c. For both correctness and KCs, we report the average human-assigned Score and two inter-rater reliability (IRR) metrics: i) the portion of times that the labels of all 3 annotators are exactly the same (Overlap) and ii) the Krippendorff’s $\boldsymbol{\alpha}$ coefficient [6]. We also compute the accuracy of correctness predictions at the final turn using CoMTA’s ground truth labels for both GPT-4o and our annotators.
@@ 7.1 Human Evaluation
Table 4 shows the human evaluation results. We see that our annotators give GPT-4o very high scores for correctness and moderate-to-high scores for KCs, indicating that GPT-4o is able to competently perform these tasks. IRR is volatile, with high overlap and low $\alpha$ on correctness, and low overlap but moderate $\alpha$ on KCs. However, the low $\alpha$ on correctness is expected given the high class imbalance of the data (more than 90% of labels are 1), and the low overlap on KCs is expected since there are 4 classes and exact ratings can be subjective. While GPT-4o’s accuracy on final turn correctness is lower than the correctness score might suggest, this is expected since the final turns are much harder to compute correctness on without subsequent tutor turns to help indicate the label. As reference, the CoMTA tech report [39] reports final turn correctness evaluation with GPT-4o at similar accuracy (78.3). Finally, given that the average human annotator’s final turn correctness accuracy is only slightly higher than GPT-4o’s, we can conclude that the correctness labeling task is a challenging one where GPT-4o achieves close to human-level performance.
@@ 7.2 Error Analysis
To understand when GPT-4o makes errors in labeling correctness and KCs, we qualitatively analyze its annotations. When labeling correctness, most of GPT-4o’s errors are on the final turn, especially in questions that require some numerical calculations to verify correctness, such as when the student provides the numerical result of a sum or product calculation. This is a known limitation in GPT-4/4o [50] and can possibly be improved through external tools such as calculators. The other common error tends to be when GPT-4o should label a turn as “na” but does not, e.g., when the tutor asks the student whether they know how to use a function on their calculator. GPT-4o also sometimes labels a turn as correct when the tutor asks a student if they know a concept and the student indicates they do not, showing that GPT-4o can be confused by the tutor’s intentions.
@@
8 Discussion, Conclusions, and Future Work
In this paper, we introduced the task of dialogueKT with the goal of estimating student knowledge from open-ended tutor-student dialogues. We used GPT-4o to automatically annotate student response correctness and relevant KCs in the dialogue turns, and showed via an expert human evaluation that these annotations are largely accurate. We introduced a new LLM-based KT method that leverages the textual content in dialogues, and showed that LLMKT i) significantly outperforms existing methods on dialogueKT and ii) works with limited training data. We also showed that our models use the textual content in dialogues to capture student knowledge in meaningful ways. There are numerous avenues for future work. First, we can jointly model tutor moves and student knowledge states to analyze what types of tutor moves are effective at improving student knowledge. Second, in addition to student knowledge, we also need to account for student affect dynamics in dialogues, possibly with the help of other data modalities such as audio and video. Third, the scope of dialogueKT can be expanded upon: in this work, we only focus on math dialogues between a single student and tutor, but future work can extend to other subjects and dialogues involving multiple students. Fourth, there are many potential applications of dialogueKT that can be explored. Future work should investigate whether estimated knowledge states are useful for personalization, providing teacher support, or can even be used to train AI-based tutors by powering simulated student agents. Overall, we hope that this initial investigation can broaden the scope of student modeling to conversational settings, which have become of significant interest lately.
9 Acknowledgments
The authors would like to thank Kristen Dicerbo and the Khan Academy team for guidelines on how to use the CoMTA dataset and releasing it. We would also like to thank Nigel Fernandez, Zichao Wang, and Jacob Whitehill for helpful discussions around this work. This work is partially supported by Schmidt Futures under the learning engineering virtual institute (LEVI) initiative and the NSF under grants 2118706, 2202506, 2237676, and 2341948.